# Finite analogs of partition bias related to hook length two and a variant of Sylvester’s map

### v1

@@ Abstract.
In this paper, we count the total number of hooks of length two in all odd partitions of $n$ and all distinct partitions of $n$ with a bound on the largest part of the partitions. We generalize inequalities of Ballantine et al by showing there is a bias in the number of hooks of length two in all odd partitions over all distinct partitions of $n$ in presence of a bound on the largest part. To establish such a bias, we use a variant of Sylvester’s map. Then, we conjecture a similar finite bias for a weighted count of hooks of length two and prove it when we remove the bound on the largest part.
@@ 1. Introduction
Clearly, $\psi$ is invertible. Consider $\psi^{-1}:=\sigma:\mathcal{D}_{n}\rightarrow\mathcal{O}_{n}$ and let $\pi_{d}\in\mathcal{D}_{n}$. Now, consider the alternating sum of parts $\gamma(\pi_{d})$ of $\pi_{d}$. We then construct a center hook consisting of $\gamma(\pi_{d})$ cells placed vertically from bottom to top to form a spine and $\lambda_{1}-\gamma(\pi_{d})$ cells placed horizontally to the right of the topmost cell of the spine. Then, symmetrically insert $\lambda_{1}-\gamma(\pi_{d})$ cells horizontally to the left of the topmost cell of the spine. Then, add $\lambda_{2}-\lambda_{1}+\gamma(\pi_{d})$ cells vertically from top to bottom below the rightmost cell of the horizontal piece to the left of the spine. Then, symmetrically insert $\lambda_{2}-\lambda_{1}+\gamma(\pi_{d})$ cells vertically from top to bottom starting from the cell in second row and in second column to the right of the spine. Now, add $\lambda_{3}-\lambda_{2}+\lambda_{1}-\gamma(\pi_{d})$ cells to the right of the cell in second row and in second column to the right of the spine. Proceeding in this fashion, after exhausting all the parts of $\pi_{d}$, we read the number of cells in the newly constructed symmetric diagram row-by-row from top to bottom to get parts of a partition $\pi_{o}$ of $n$ into odd parts, i.e., $\pi_{o}\in\mathcal{O}_{n}$.
@@
(1.3) $\sum\limits_{n\geq 0}a_{2}(n)q^{n}=\dfrac{1}{(q^{2};q^{2})_{\infty}}\left(q^{2}+\sum\limits_{n\geq 2}\left(q^{2n-1}+q^{2(2n-1)}\right)\right).$
@@ 3. Proofs of Main Results
To prove Proposition 3.1, we need the following lemma due to Alladi and Berkovich [1, Lemma $3$].
@@
Let $h_{L,i}(y,q)$ denote the generating function of unrestricted partitions $\pi$ having $l(\pi)\leq L$ into exactly $i$ non-negative parts such that $\pi$ is counted with weight $(1-y)^{\nu^{+}_{d}(\pi)}$, where $\nu^{+}_{d}(\pi)$ is the number of different positive parts of $\pi$. Then
@@
We now illustrate $\varphi$ with two examples:
@@ Acknowledgments
We would like to thank Krishnaswami Alladi, George E. Andrews, Ali K. Uncu and Hamza Yesilyurt for their kind interest.

### v2

@@ Abstract.
In this paper, we count the total number of hooks of length two in all odd partitions of $n$ and all distinct partitions of $n$ with a bound on the largest part of the partitions. We generalize inequalities of Ballantine, Burson, Craig, Folsom and Wen by showing there is a bias in the number of hooks of length two in all odd partitions over all distinct partitions of $n$ in presence of a bound on the largest part. To establish such a bias, we use a variant of Sylvester’s map. Then, we conjecture a similar finite bias for a weighted count of hooks of length two and prove it when we remove the bound on the largest part.
@@ 1. Introduction
Clearly, $\psi$ is invertible. Consider $\psi^{-1}:=\sigma:\mathcal{D}_{n}\rightarrow\mathcal{O}_{n}$ and let $\pi_{d}=(\lambda_{1},\lambda_{2},\lambda_{3},\ldots)\in\mathcal{D}_{n}$. Now, consider the alternating sum of parts $\gamma(\pi_{d})$ of $\pi_{d}$. We then construct a center hook consisting of $\gamma(\pi_{d})$ cells placed vertically from bottom to top to form a spine and $\lambda_{1}-\gamma(\pi_{d})$ cells placed horizontally to the right of the topmost cell of the spine. Then, symmetrically insert $\lambda_{1}-\gamma(\pi_{d})$ cells horizontally to the left of the topmost cell of the spine. Then, add $\lambda_{2}-\lambda_{1}+\gamma(\pi_{d})$ cells vertically from top to bottom below the rightmost cell of the horizontal piece to the left of the spine. Then, symmetrically insert $\lambda_{2}-\lambda_{1}+\gamma(\pi_{d})$ cells vertically from top to bottom starting from the cell in second row and in second column to the right of the spine. Now, add $\lambda_{3}-\lambda_{2}+\lambda_{1}-\gamma(\pi_{d})$ cells to the right of the cell in second row and in second column to the right of the spine. Proceeding in this fashion, after exhausting all the parts of $\pi_{d}$, we read the number of cells in the newly constructed symmetric diagram row-by-row from top to bottom to get parts of a partition $\pi_{o}$ of $n$ into odd parts, i.e., $\pi_{o}\in\mathcal{O}_{n}$.
@@
Conjecture 1.4 was recently proved by Craig, Dawsey and Han (see [6, Theorem $1.2$]).
@@
(1.3) $\sum\limits_{n\geq 0}a_{2}(n)q^{n}=\dfrac{1}{(q;q^{2})_{\infty}}\left(q^{2}+\sum\limits_{n\geq 2}\left(q^{2n-1}+q^{2(2n-1)}\right)\right).$
@@ 3. Proofs of Main Results
To prove Proposition 3.1, we need the following lemma.
@@
(Alladi-Berkovich [1, Lemma $3$]) Let $h_{L,i}(y,q)$ denote the generating function of unrestricted partitions $\pi$ having $l(\pi)\leq L$ into exactly $i$ non-negative parts such that $\pi$ is counted with weight $(1-y)^{\nu^{+}_{d}(\pi)}$, where $\nu^{+}_{d}(\pi)$ is the number of different positive parts of $\pi$. Then
@@
We now illustrate $\varphi$ with three examples:
Example 3.4.
Let $L=7$ and $n=11$. Consider $(\pi_{1},\pi_{2})=((\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}),\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\emptyset\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})\in\mathcal{B}_{7,11}$. Then, by Case I, we have
$\varphi((\pi_{1},\pi_{2}))$ $=\varphi(((\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}),\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\emptyset\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}))$
$=(\sigma((\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})),\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\emptyset\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$
=(($11\displaystyle=((\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{11}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}),\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\emptyset\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})\in\mathcal{A}_{7,11}.$
[…]
Figure 5. $\varphi(((\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}),\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\emptyset\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}))=((\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}),\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\emptyset\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$
@@ Acknowledgments
We would like to thank Krishnaswami Alladi, George E. Andrews, Ali K. Uncu and Hamza Yesilyurt for their kind interest. We also thank Will Craig for directing us to [6] for a proof of Conjecture 1.4.